OpenStudy (anonymous):

log(x+5)+log(5)=3 What's x?

OpenStudy (anonymous):

Or rather, how do i solve for x?

OpenStudy (anonymous):

\[\log a+\log b=\log ab\]

OpenStudy (anonymous):

\[\log(5(x+5))=3\]

OpenStudy (anonymous):

Um.. the answer's actually not 10

OpenStudy (anonymous):

At least that's what my calculator is telling me

OpenStudy (anonymous):

\[5x+25=10^3\]